OpenStudy (anonymous):

Is this a polynomial function? if so, write in standard form. f(x)=5x+2

OpenStudy (diyadiya):

its a linear polynomial , and it is in the standard form

OpenStudy (anonymous):

it is and it is

OpenStudy (anonymous):

okay, what about: f(x)=2^x? I dont think its a polynomial

OpenStudy (diyadiya):

Any linear polynomial is of the form ax + b, where x is a variable and a, b are constants

OpenStudy (jamesj):

@green: what's the definition of a polynomial?

OpenStudy (anonymous):

a monomial of sum of monomials.

OpenStudy (jamesj):

And what is a monomials?

OpenStudy (jamesj):

a monomial?

OpenStudy (anonymous):

#, variable, or a product of them (ex. 5x)

OpenStudy (jamesj):

Right. So is 2^x a monomial or the sum of them?

OpenStudy (jamesj):

(also monomials can be terms like x^3, 5x^4, -20000x^1000000)

OpenStudy (anonymous):

okay, um, h(x0=1/2x^4-x^2+3x^3 in stadard form is h(x)=1/2x^4+3x^3-x^2?

OpenStudy (jamesj):

is h(x) the sum of monomials?

OpenStudy (anonymous):

I see the mistake, sorry. *h(x)

OpenStudy (jamesj):

No, it's not. So h(x) is not a polynomial.

OpenStudy (anonymous):

wait, but how is it different from f(x) or g(x)?